# LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning

### v2

@@ A.1 Implementation details
A.2 Environment details
In this subsection, we provide the environment and implementation details for all the tasks:
@@ A.2 Environment details
Pick and place and bin environments In this subsection, we describe the environment details for the pick and place and bin tasks. The state is represented as the vector $[dx,o,q,e]$, where $dx$ is the current gripper position, $o$ is the position of the block object on the table, $q$ is the relative position of the block with respect to the gripper, and $e$ includes the linear and angular velocities of both the gripper and the block object. The higher-level policy input is a concatenated vector $[dx,o,q,e,g]$, where $g$ is the target goal position. The lower-level policy input is a concatenated vector $[dx,o,q,e,s_{g}]$, where $s_{g}$ is the sub-goal provided by the higher-level policy. In our experiments, the sizes of $dx$, $o$, $q$, and $e$ are set to $3$, $3$, $3$, and $11$, respectively. The lower primitive action $a$ is a 4-dimensional vector with each dimension $a_{i}\in[0,1]$. The first three dimensions provide gripper position offsets, and the last dimension controls the gripper. During training, the positions of the block object and the goal are randomly generated (the block is always initialized on the table, and the goal is always above the table at a fixed height).
Franka kitchen environment For this environment please refer to the D4RL environment [76]. In this environment, the franka robot has to perform a complex multi-stage task in order to achieve the final goal.
A.3 Full Prompts
@@
[backgroundcolor=lightgray] We want you to generate a random position for an object within the table following the description and rules.
@@
#generate maze with all the constarints
@@
Motion Descriptor Prompt for Pick and Place Environment {mdframed}[backgroundcolor=lightgray] We want you to generate a random position for an object within the table following the description and rules.
@@
Reward Generator Prompt for Pick and Place Environment {mdframed}[backgroundcolor=lightgray]
@@
Motion descriptor prompt for Bin environment {mdframed}[backgroundcolor=lightgray] We want you to generate a random position for a bin and an object within the bin following the description and rules. [Description]
@@
Reward generator prompt for Bin environment {mdframed}[backgroundcolor=lightgray] We have a description of a robot’s motion and we want you to turn that into the corresponding program with following functions:
@@
[backgroundcolor=lightgray] We want you to generate a random position for the door of a microwave and gas-knob following the description and rules.
@@
Reward generator prompt for Franka kitchen environment We have a description of a robot’s motion and we want you to turn that into the corresponding program with following functions:
@@
A.4 Sample Codes
In this section, we show sampled responses for all four environments mentioned in this paper for a better understanding of the LGR2 method. We sample 2-3 responses for each method to show both successful and failed results. Sample Code for Maze Navigation environment This section provides the sample code for the Maze Navigation environment generated by GPT-3.5. {mdframed}[backgroundcolor=lightgray]
@@ A.4 Sample Codes
print(”Environment reset.”)
@@
print(f”Setting gripper position to x: x_pos, y: y_pos, z: z_pos.”)
@@
print(”Generating maze…”)
@@
print(”Running simulation…”)
@@
This section provides the sample code for the Pick and Place environment generated by GPT-3.5.
@@
print(”Environment reset.”)
@@
print(f”Setting gripper position to x: x_pos, y: y_pos, z: z_pos.”)
def generate_Object_Pos(): print(”Generating object position.”)
@@
print(”Simulation started.”)
@@
This section provides the sample code for the Bin environment generated by GPT-3.5.
@@
print(”Environment reset.”)
@@
print(f”Setting gripper position to x: x_pos, y: y_pos, z: z_pos.”)
@@
print(”Generating bin with all constraints.”)
@@
print(”Simulation started.”)
@@
This section provides the sample code for the Franka kitchen environment generated by GPT-3.5.
@@
print(f”Setting microwave door position to joint_1_Pos.”)
@@
print(f”Setting gas knob position to x: joint_1_Pos, y: joint_2_Pos.”)
@@
print(”Simulation started.”)
@@
A.5 Impact Statement
Our proposed approach and algorithm are not expected to bring about immediate technological breakthroughs. Instead, our primary contributions are conceptual, emphasizing fundamental aspects of Hierarchical Reinforcement Learning (HRL). This foundational work sets the stage for future research and could foster advancements in HRL and its associated areas.
A.6 Qualitative visualizations

### v3

@@ Appendix B Implementation details
Appendix C Environment details
In this section, we provide the environment and implementation details for all the tasks:
@@ Appendix C Environment details
Pick and place and bin environments In this section, we describe the environment details for the pick and place and bin tasks. The state is represented as the vector $[dx,o,q,e]$, where $dx$ is the current gripper position, $o$ is the position of the block object on the table, $q$ is the relative position of the block with respect to the gripper, and $e$ includes the linear and angular velocities of both the gripper and the block object. The higher-level policy input is a concatenated vector $[dx,o,q,e,g]$, where $g$ is the target goal position. The lower-level policy input is a concatenated vector $[dx,o,q,e,s_{g}]$, where $s_{g}$ is the sub-goal provided by the higher-level policy. In our experiments, the sizes of $dx$, $o$, $q$, and $e$ are set to $3$, $3$, $3$, and $11$, respectively. The lower primitive action $a$ is a 4-dimensional vector with each dimension $a_{i}\in[0,1]$. The first three dimensions provide gripper position offsets, and the last dimension controls the gripper. During training, the positions of the block object and the goal are randomly generated (the block is always initialized on the table, and the goal is always above the table at a fixed height).
Franka kitchen environment For this environment please refer to the D4RL environment Fu et al. (2020). In this environment, the franka robot has to perform a complex multi-stage task in order to achieve the final goal.
Appendix D Full Prompts
@@
[backgroundcolor=white] We want you to generate a random position for an object within the table following the description and rules.
@@
#generate maze with all the constraints
@@
Motion Descriptor Prompt for Pick and Place Environment {mdframed}[backgroundcolor=white] We want you to generate a random position for an object within the table following the description and rules.
@@
Reward Generator Prompt for Pick and Place Environment {mdframed}[backgroundcolor=white]
@@
Motion descriptor prompt for Bin environment {mdframed}[backgroundcolor=white] We want you to generate a random position for a bin and an object within the bin following the description and rules. [Description]
@@
Reward generator prompt for Bin environment {mdframed}[backgroundcolor=white] We have a description of a robot’s motion and we want you to turn that into the corresponding program with following functions:
@@
[backgroundcolor=white] We want you to generate a random position for the door of a microwave and gas-knob following the description and rules.
@@
Reward generator prompt for Franka kitchen environment
@@
Appendix E Sample Codes
In this section, we show sampled responses for all four environments mentioned in this paper for a better understanding of the LGR2 method. We sample 2-3 responses for each method to show both successful and failed results.
Sample Code for Maze Navigation environment Here, we provide the sample code for the Maze Navigation environment generated by GPT-3.5. {mdframed}[backgroundcolor=white]
@@ Appendix E Sample Codes
print("Environment reset.")
@@
print(f"Setting gripper position to x: x_pos, y: y_pos, z: z_pos.")
@@
print("Generating maze…")
@@
print("Running simulation…")
@@
Here, we provide the sample code for the Pick and Place environment generated by GPT-3.5.
@@
print("Environment reset.")
@@
print(f"Setting gripper position to x: x_pos, y: y_pos, z: z_pos.")
def generate_Object_Pos(): print("Generating object position.")
@@
print("Simulation started.")
@@
Here, we provide the sample code for the Bin environment generated by GPT-3.5.
@@
print("Environment reset.")
@@
print(f"Setting gripper position to x: x_pos, y: y_pos, z: z_pos.")
@@
print("Generating bin with all constraints.")
@@
print("Simulation started.")
@@
Here, we provide the sample code for the Franka kitchen environment generated by GPT-3.5.
@@
print(f"Setting microwave door position to joint_1_Pos.")
@@
print(f"Setting gas knob position to x: joint_1_Pos, y: joint_2_Pos.")
@@
print("Simulation started.")
@@
Appendix F Qualitative visualizations